# Instance-Optimal Imprecise Convex Hull

### v3

@@ Abstract
We construct the convex hull of an imprecise point set in the plane, by determining the cyclic ordering of the convex hull vertices of $P$ as efficiently as possible. Efficiency is interpreted in two ways: (i) minimising the number of retrievals, and (ii) computing each region to retrieve quickly.
@@
Compared to previous approaches that achieve optimal retrieval counts, we improve the runtime per retrieval from polynomial to polylogarithmic. We extend our results to simple $k$-gons, to pairwise disjoint disks with radii in $[1,k]$, and to unit disks where at most $k$ disks overlap in a single point. Our runtime scales linearly with $k$.
@@ 1 Introduction
Imprecision is inherent in real-world data. It arises from rounding errors in floating-point computations, inaccuracies in measurement, and delayed sampling in GPS devices. In many scenarios, imprecise data can be refined at a cost by computing exact values or by taking additional samples. This is formalised in the model of imprecise geometry introduced by Held and Mitchell [14], and studied further in [9, 14, 25, 10, 11, 19, 17].
@@
An imprecise point set [14] is defined as a family of regions $F=(R_{1},\ldots,R_{n})$, where each region $R_{i}$ contains a unique but unknown point $p_{i}$. A realisation $P\sim F$ is a sequence $P=(p_{1},\ldots,p_{n})$ where $p_{i}\in R_{i}$. An input instance is a pair $(F,P)$. A retrieval operation reveals the precise location of a point $p_{i}$, replacing $R_{i}$ with $p_{i}$.
[…]
@@
Many geometric problems have been studied in imprecise geometry, including Delaunay triangulations [9, 14, 19, 25], convex hulls [10, 11], Gabriel graphs [19], and onion decompositions [17]. For these geometric problems, it is known that any reconstruction algorithm must, in the worst case, use $\Omega(n)$ retrievals. Most of the existing algorithms assume that the regions in $F$ are pairwise disjoint unit disks [9, 14, 19, 25, 10, 17]. Under these assumptions, deterministic worst-case optimal algorithms exist with $O(n\log n)$ preprocessing time, $O(n)$ retrievals, and $O(n)$ reconstruction time. Ezra and Mulzer [11] allow $F$ to consist of arbitrary lines reconstruct the convex hull using $O(n)$ retrievals and $O(n\cdot\alpha(n))$ expected time where $\alpha(n)$ denotes the inverse Ackermann function. Buchin, Löffler, Morin and Mulzer [4] permit overlapping unit disks of ply $k$ (i.e., at most $k$ disks intersect each point in the plane). We summarize these results in the top three rows of Table 1.
@@
Afshani, Barbay, and Chan [1] proved that, for many geometric problems including constructing the convex hull, instance-optimal reconstruction time is unachievable. Likewise, it is easy to see that instance-optimality in the number of retrievals cannot be guaranteed in general: E.g., let $F$ consist of two overlapping intervals $R_{1}$ and $R_{2}$ in $\mathbb{R}$, with $R_{1}\setminus R_{2}\neq\emptyset$ and $R_{2}\setminus R_{1}\neq\emptyset$. Define $P_{1}$ with $p_{1}\in R_{1}\setminus R_{2}$ and $p_{2}\in R_{1}\cap R_{2}$, and $P_{2}$ with $p_{1}\in R_{1}\cap R_{2}$ and $p_{2}\in R_{2}\setminus R_{1}$. In $(F,P_{1})$, retrieving $R_{1}$ suffices; in $(F,P_{2})$, retrieving $R_{2}$ suffices. No algorithm can distinguish between the two instances without making a retrieval, and hence must use two retrievals on one of the instances. Since exact instance-optimality cannot be obtained, we instead aim for asymptotically instance-optimality.
[…]
@@
Only a few instance-optimal reconstruction algorithms are known (see Table 1). Bruce, Hoffmann, Krizanc, and Raman [3] presented the first such algorithm for the convex hull. Their approach is computationally expensive, using an unspecified but superlinear time per retrieval (we discuss this in detail in Appendix A). Subsequent works show that polylogarithmic time per retrieval is achievable, albeit on geometric structures that are much simpler than the convex hull.
Van der Hoog, Kostitsyna, Löffler, and Speckmann [21] introduced the first near-linear instance-optimal algorithm, solving the sorting problem for one-dimensional intervals. Their algorithm preprocesses the intervals in $O(n\log n)$ time and uses $\Theta(r(F,P))$ retrievals, with at most logarithmic time per retrieval. Later, Van der Hoog, Kostitsyna, Löffler, and Speckmann [22] extended this approach to two-dimensional inputs for Pareto front reconstruction, under the assumption that $F$ consists of pairwise disjoint axis-aligned rectangles. Their method also guarantees at most logarithmic time per retrieval.
@@
Simultaneously and independently of this paper, Löffler and Raichel [18] developed an algorithm for reconstructing the convex hull when $F$ is a set of unit disks of ply $k$. Their number of retrievals (Table 1) is not instance-optimal but rather worst-case optimal for every instance $F$. Formally, they use $O(w(F))$ retrievals for some region-dependent value $w(F)$ with $r(F,P)\leq w(F)\leq n$. We discuss their result in more detail in Appendix A.
@@
The study of instance-optimal algorithms extends beyond computational geometry. A prime example is sorting under partial information. Given a partial order $O$ over a set $X$ and an unknown linear extension $L$, the goal is to sort $X$ using a minimum number of comparisons, where a comparison queries the order of a pair $(x,y)\in X$ under $L$. This is directly analogous to retrievals. Kahn and Saks [16] introduced an exponential-time algorithm that is instance-optimal in the number of comparisons. Since then, significant progress has been made on improving the runtime of such algorithms [15, 6], culminating in near-linear time algorithms that are instance-optimal in both runtime and comparisons [24, 23, 13]. Another example is the bidirectional shortest path problem. Haeupler, Hladík, Rozhoň, Tarjan, and Tětek [12] show an algorithm that finds the shortest path between two nodes $s$ and $t$ using an instance-optimal number of edge-weight comparisons.
@@
We present the first algorithm for convex hull reconstruction that is instance-optimal while only requiring polylogarithmic time per retrieval. If $F$ is a family of $n$ constant-complexity simple polygons, we preprocess $F$ in $O(n\log^{3}n)$ time and reconstruct the convex hull of any $P\sim F$ using $\Theta(r(F,P))$ retrievals and $O(r(F,P)\log^{3}n)$ total time. Our approach generalises to simple $k$-gons with a linear factor in $k$ in space and time, to pairwise disjoint disks with radii in $[1,k]$, and to unit disks of ply $k$ (see Table 1). Compared to previous approaches for convex hulls that achieve optimal retrieval counts [3], we exponentially improve the runtime per retrieval from polynomial to polylogarithmic. Compared to near-linear time algorithms for convex hulls (or, Delaunay triangulations) [9, 14, 19, 18, 25, 10, 11], we significantly reduce the number of retrievals used and broaden the generality of input families. The latter comes at a cost of a polylogarithmic slowdown if $r(F,P)$ is linear.
@@
Unit disks No many $O(n\log n)$ $O(n)$ $O(n)$ [9, 10, 14, 17, 19, 25]
Disks $k$ Delaunay $O(n\log n)$ $O(n)$ $O(n\log k)$ rand. [5]
Lines Yes hull $O(n\log n)$ $O(n)$ $(n\cdot\alpha(n))$ rand. [11]
Intervals Yes sorting $O(n\log n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\log n)$ [21]
Axis rect. No front $O(n\log n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\log n)$ [22]
Smooth Yes front, hull $O(\operatorname{poly}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\operatorname{poly}n)$ [3]
@@
$[1,k]$-disks No hull $O(kn\log^{3}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot k\log^{3}n)$ Thm 6.31
Unit disks $k$ hull $O(kn\log^{3}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot k\log^{3}n)$ Thm 6.31
Unit disks $k$ hull $O(k^{3}n)$ $O(w(F))$ $O(k^{3}\cdot w(F))$ [18]
@@
In Section 2, we provide some preliminaries. Then, in Section 3, we present our general algorithm for reconstructing the convex hull and prove its optimality (Theorem 3.18). In Section 4, we present a data structure for simple $k$-gons and give an instance-optimal algorithm that uses $O(kn)$ space and $O(r(F,P)\cdot k^{2}\log^{3}(kn))$ time (Theorem 4.17). In Section 5 we improve the dependency on $k$ to near-linear (Theorem 4.19). In Section 6, we study pairwise disjoint disks with a radius in $[1,k]$, or, unit disks with ply $k$.
@@ 2 Preliminaries
Recall that an imprecise point set is defined as a family of geometric regions $F=(R_{1},\ldots,R_{n})$, and a realisation $P\sim F$ is defined as a sequence $P$ of $n$ points with $p_{i}\in R_{i}$. We explicitly allow for duplicate points in $P$ and do not assume general position. The algorithm has two phases [14]: a preprocessing phase, where only $F$ is available, and a reconstruction phase, in which we may perform a retrieval on any $R_{i}\in F$, replacing $R_{i}$ with the point $p_{i}$. To maintain generality, we require only that each $R_{i}$ is a point or a closed, connected, bounded region whose boundary is a simple closed piecewise-$C^{1}$ curve. Non-point regions must have connected interiors and coincide with the closure of their interiors; see Figure 2. Next, we define retrievals, convex hulls, vertices, edges, reconstruction strategies, and instance-optimality.
@@
[…]
@@
Since all regions in $F$ are closed, we must assume $P$ may contain duplicates and require that the convex hull of a point sequence includes all collinear and duplicate points. To illustrate, let $F$ consist of $n$ identical unit squares, and suppose $P\sim F$ includes four points forming the corners of one square. A naïve algorithm might retrieve these four points first and construct the convex hull. If we assumed general position or excluded duplicates from the convex hull, this algorithm could then conclude that all remaining points cannot lie on the convex hull and terminate early. However, such behaviour cannot be guaranteed against an adversary. Since all regions are indistinguishable, an adversary may simply give these four points last. Alternatively, if $F$ consisted of open regions, we could assume general position and exclude duplicates.
@@
[…]
@@
[…]
A reconstruction algorithm is instance-optimal if it retrieves $\Theta(r(F,P))$ regions before $F$ is finished. We distinguish two types of reconstruction algorithms:
@@
Bruce, Hoffmann, Krizanc, and Raman [3] present an instance-optimal reconstruction strategy for region families $F$ consisting of points and closed piecewise-$C^{1}$ regions. They do not present a corresponding reconstruction program. When $F$ consists of disks or polygons, a reconstruction program may be derived, albeit with unspecified polynomial-time costs per retrieval. We further discuss their algorithm in Appendix A. We rely on a key concept of their paper, i.e. a witness.

### v4

@@ Abstract
We construct the convex hull of an imprecise point set in the plane, by determining the cyclic ordering of the convex hull vertices of $P$ as efficiently as possible. Efficiency is interpreted in two ways: (i) minimising the number of retrievals, and (ii) the computation time to determine the set of regions that must be retrieved.
@@
Compared to previous approaches that achieve optimal retrieval counts, we improve the runtime per retrieval from polynomial to polylogarithmic. We extend the generality of previous results to simple $k$-gons, to pairwise disjoint disks with radii in $[1,k]$, and to unit disks where at most $k$ disks overlap in a single point. Our runtime scales linearly with $k$.
@@ 1 Introduction
Imprecision is inherent in real-world data. It arises from rounding errors in floating-point computations, inaccuracies in measurement, and delayed sampling in GPS devices. In many scenarios, imprecise data can be refined at a cost by computing exact values or by taking additional samples. This is formalised in the model of imprecise geometry introduced by Held and Mitchell [held2008triangulating], and studied further in [devillers2011delaunay, held2008triangulating, van2010preprocessing, evans2011possible, ezra2013convex, loffler2010delaunay, loffler2013unions].
@@
An imprecise point set [held2008triangulating] is defined as a family of regions $F=(R_{1},\ldots,R_{n})$, where each region $R_{i}$ contains a unique but unknown point $p_{i}$. A realisation $P\sim F$ is a sequence $P=(p_{1},\ldots,p_{n})$ where $p_{i}\in R_{i}$. An input instance is a pair $(F,P)$. A retrieval operation reveals the precise location of a point $p_{i}$, replacing $R_{i}$ with $p_{i}$.
Let $F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptB$ denote the family $F$ after retrieving all $R_{i}\in B$, where $B\subset F$. The aim of a reconstruction algorithm is to identify a subset $B\subset F$ such that for all realisations $P_{1},P_{2}\sim F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptB$, the algorithm’s output (e.g., the cyclic order of the region indices of the points around the convex hull) is identical for $P_{1}$ and $P_{2}$. Figure 1 illustrates an example. After retrieving the subset $B=\{R_{3},R_{4}\}$, the cycling ordering of the convex hull vertices is identically $(p_{1},p_{3},p_{2},p_{4})$ for all realisations of $F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptB$.
@@
Many geometric problems have been studied in imprecise geometry, including Delaunay triangulations [devillers2011delaunay, held2008triangulating, loffler2010delaunay, van2010preprocessing], convex hulls [evans2011possible, ezra2013convex], Gabriel graphs [loffler2010delaunay], and onion decompositions [loffler2013unions]. For these geometric problems, it is known that any reconstruction algorithm must, in the worst case, use $\Omega(n)$ retrievals. Most of the existing algorithms assume that the regions in $F$ are pairwise disjoint unit disks [devillers2011delaunay, held2008triangulating, loffler2010delaunay, van2010preprocessing, evans2011possible, loffler2013unions]. Under these assumptions, deterministic worst-case optimal algorithms exist with $O(n\log n)$ preprocessing time, $O(n)$ retrievals, and $O(n)$ reconstruction time. Ezra and Mulzer [ezra2013convex] allow $F$ to consist of arbitrary lines and reconstruct the convex hull using $O(n)$ retrievals and $O(n\cdot\alpha(n))$ expected time where $\alpha(n)$ denotes the inverse Ackermann function. Buchin, Löffler, Morin and Mulzer [buchin2011preprocessing] permit overlapping unit disks of ply $k$ (i.e., at most $k$ disks intersect each point in the plane). We summarize these results in the top three rows of Table 1.
@@
Afshani, Barbay, and Chan [afshani2017instance] proved that, for many geometric problems including constructing the convex hull, instance-optimal reconstruction time is unachievable. Likewise, it is easy to see that instance-optimality in the number of retrievals cannot be guaranteed in general: E.g., let $F$ consist of two overlapping intervals $R_{1}$ and $R_{2}$ in $\mathbb{R}$, with $R_{1}\setminus R_{2}\neq\emptyset$ and $R_{2}\setminus R_{1}\neq\emptyset$. Define $P_{1}$ with $p_{1}\in R_{1}\setminus R_{2}$ and $p_{2}\in R_{1}\cap R_{2}$, and $P_{2}$ with $p_{1}\in R_{1}\cap R_{2}$ and $p_{2}\in R_{2}\setminus R_{1}$. In $(F,P_{1})$, retrieving $R_{1}$ suffices; in $(F,P_{2})$, retrieving $R_{2}$ suffices. No algorithm can distinguish between the two instances without making a retrieval, and hence must use two retrievals on one of the instances. Since exact instance-optimality cannot be obtained, we instead aim for asymptotic instance-optimality.
Formally, we denote for any input instance by $r(F,P)$ the minimum integer such that there exists a subset $B\subset F$ of size $r(F,P)$ where all realizations $P^{\prime}\sim(F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptB)$ have the same vertex-ordering along the convex hull of $P^{\prime}$. We note that $r(F,P)$ is, equivalently, the optimal number of retrievals. We say that an algorithm is instance-optimal if for all inputs $(F,P)$ it uses $\Theta(r(F,P))$ retrievals. We highlight that $r(F,P)$ depends on both the region family $F$ and the specific realisation $P$. To illustrate, suppose $F$ consists of nested rectangles $R_{n}\subset R_{n-1}\subset\ldots\subset R_{1}$. If $P\sim F$ is such that the convex hull of $(p_{1},p_{2},p_{3},p_{4})$ contains all of $R_{5}$, then retrieving just those four suffices: $r(F,P)=4$. But if all $p_{i}$ coincide then $r(F,P)=n$.
@@
Only a few instance-optimal reconstruction algorithms are known (see Table 1). Bruce, Hoffmann, Krizanc, and Raman [bruce2005efficient] presented the first such algorithm for the convex hull. Their approach is computationally expensive, using an unspecified but superlinear time per retrieval (we discuss this in the appendix). Subsequent works show that polylogarithmic time per retrieval is achievable, albeit on geometric structures that are much simpler than the convex hull.
Van der Hoog, Kostitsyna, Löffler, and Speckmann [van2019preprocessing] introduced the first near-linear instance-optimal algorithm, solving the sorting problem for one-dimensional intervals. Their algorithm preprocesses the intervals in $O(n\log n)$ time and uses $\Theta(r(F,P))$ retrievals, with at most logarithmic time per retrieval. Later, Van der Hoog, Kostitsyna, Löffler, and Speckmann [van2022preprocessing] extended this approach to two-dimensional inputs for Pareto front reconstruction, under the assumption that $F$ consists of pairwise disjoint axis-aligned rectangles. Their method also guarantees at most logarithmic time per retrieval.
@@
Simultaneously and independently of this paper, Löffler and Raichel [loffler2025preprocessing] developed an algorithm for reconstructing the convex hull when $F$ is a set of unit disks of ply $k$. Their number of retrievals (Table 1) is not instance-optimal but rather worst-case optimal for every instance $F$. Formally, they use $O(w(F))$ retrievals for some region-dependent value $w(F)$ with $r(F,P)\leq w(F)\leq n$. We discuss their result in the appendix.
@@
The study of instance-optimal algorithms extends beyond computational geometry. A prime example is sorting under partial information. Given a partial order $O$ over a set $X$ and an unknown linear extension $L$, the goal is to sort $X$ using a minimum number of comparisons, where a comparison queries the order of a pair $(x,y)\in X$ under $L$. This is directly analogous to retrievals. Kahn and Saks [kahn_balancing_1984] introduced an exponential-time algorithm that is instance-optimal in the number of comparisons. Since then, significant progress has been made on improving the runtime of such algorithms [kahn_entropy_1992, cardinal_sorting_2013], culminating in near-linear time algorithms that are instance-optimal in both runtime and comparisons [van2024tight, van2025simpler, Haeupler25]. Another example is the bidirectional shortest path problem. Haeupler, Hladík, Rozhoň, Tarjan, and Tětek [haeupler2025bidirectional] show an algorithm that finds the shortest path between two nodes $s$ and $t$ using an instance-optimal number of edge-weight comparisons.
@@
We present the first algorithm for convex hull reconstruction that is instance-optimal while only requiring polylogarithmic time per retrieval. If $F$ is a family of $n$ constant-complexity simple polygons, we preprocess $F$ in $O(n\log^{3}n)$ time and reconstruct the convex hull of any $P\sim F$ using $\Theta(r(F,P))$ retrievals and $O(r(F,P)\log^{3}n)$ total time. Our approach applies to simple $k$-gons with a linear factor in $k$ in space and time, to pairwise disjoint disks with radii in $[1,k]$, and to unit disks of ply $k$ (see Table 1). Compared to previous approaches for convex hulls that achieve optimal retrieval counts [bruce2005efficient], we exponentially improve the runtime per retrieval from polynomial to polylogarithmic. Compared to near-linear time algorithms for convex hulls (or, Delaunay triangulations) [devillers2011delaunay, held2008triangulating, loffler2010delaunay, loffler2025preprocessing, van2010preprocessing, evans2011possible, ezra2013convex], we significantly reduce the number of retrievals used and broaden the generality of input families. The latter comes at a cost of a polylogarithmic slowdown if $r(F,P)$ is linear in $n$.
@@
Unit disks No many $O(n\log n)$ $O(n)$ $O(n)$ [devillers2011delaunay, evans2011possible, held2008triangulating, loffler2013unions, loffler2010delaunay, van2010preprocessing]
Disks $k$ Delaunay $O(n\log n)$ $O(n)$ $O(n\log k)$ rand. [buchin2011delaunay]
Lines Yes hull $O(n\log n)$ $O(n)$ $(n\cdot\alpha(n))$ rand. [ezra2013convex]
Intervals Yes sorting $O(n\log n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\log n)$ [van2019preprocessing]
Axis rect. No front $O(n\log n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\log n)$ [van2022preprocessing]
Smooth Yes front, hull $O(\operatorname{poly}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot\operatorname{poly}n)$ [bruce2005efficient]
@@
$[1,k]$-disks No hull $O(kn\log^{3}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot k\log^{3}n)$ Thm. 6.31
Unit disks $k$ hull $O(kn\log^{3}n)$ $\Theta(r(F,P))$ $O(r(F,P)\cdot k\log^{3}n)$ Thm. 6.31
Unit disks $k$ hull $O(k^{3}n)$ $O(w(F))$ $O(k^{3}\cdot w(F))$ [loffler2025preprocessing]
@@
In Section 2, we provide some preliminaries. Then, in Section 3, we present our general algorithm for reconstructing the convex hull and prove its optimality (Theorem˜3.18). In Section 4, we present a data structure for simple $k$-gons and give an instance-optimal algorithm that uses $O(kn)$ space and $O(r(F,P)\cdot k^{2}\log^{3}(kn))$ time (Theorem 4.17). In the full version we improve the dependency on $k$ to near-linear (Theorem 4.19). In the full version, we study pairwise disjoint disks with a radius in $[1,k]$, or, unit disks with ply $k$.
@@ 2 Preliminaries
Recall that an imprecise point set is defined as a family of geometric regions $F=(R_{1},\ldots,R_{n})$, and a realisation $P\sim F$ is defined as a sequence $P$ of $n$ points with $p_{i}\in R_{i}$. We explicitly allow for duplicate points in $P$ and do not assume general position. The algorithm has two phases [held2008triangulating]: a preprocessing phase, where only $F$ is available, and a reconstruction phase, in which we may perform a retrieval on any $R_{i}\in F$, replacing $R_{i}$ with the point $p_{i}$. To maintain generality, we require only that each $R_{i}$ is a point or a closed, connected, bounded region whose boundary is a simple closed piecewise-$C^{1}$ curve. Non-point regions must have connected interiors and coincide with the closure of their interiors; see Figure 2. Next, we define retrievals, convex hulls, vertices, edges, reconstruction strategies, and instance-optimality.
@@
Given a family of regions $F$, a retrieval operation on a non-point region $R_{i}$ replaces $R_{i}$ with its associated point $p_{i}$ (thereby updating $F$). We write $F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptA$ for the family resulting from retrieving all regions in $A\subseteq F$. We denote by FAF-A the family obtained by removing all regions in $A$ from $F$. This removal is used in our analysis when we identify regions whose corresponding points cannot appear on the convex hull.
@@
Since all regions in $F$ are closed, we claim that we must assume $P$ may contain duplicates and require that the convex hull of a point sequence includes all collinear and duplicate points. To illustrate, let $F$ consist of $n$ identical unit squares, and suppose $P\sim F$ includes four points forming the corners of one square. A lucky algorithm might retrieve these four points first and construct the convex hull [Kahan1991Model]. If we assumed general position or excluded duplicates from the convex hull, this algorithm could then conclude that all remaining points cannot lie on the convex hull and terminate early. However, such behaviour cannot be guaranteed against an adversary. Since all regions are indistinguishable, an adversary may simply give these four points last. Alternatively, if $F$ consisted of open regions, we could assume general position and exclude duplicates.
@@
We say a family $F$ is finished if all realisations $P_{1},P_{2}\sim F$ satisfy $\operatorname{\preceq}(\operatorname{CH}(P_{1}))=\operatorname{\preceq}(\operatorname{CH}(P_{2}))$. A reconstruction algorithm retrieves some $B\subseteq F$ such that $F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptB$ is finished. Let $r(F,P)$ denote the minimum number of retrievals needed by any such algorithm.
@@
For any fixed $P\sim F$, let $A\subset F$ be a smallest subset of $F$ such that $F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptA$ is finished. Then $|A|$ is a tight lower bound for $r(F,P)$.
A reconstruction algorithm is instance-optimal if for all inputs $(F,P)$ it retrieves $\Theta(r(F,P))$ regions before $F$ is finished. We distinguish two types of reconstruction algorithms:
@@
Bruce, Hoffmann, Krizanc, and Raman [bruce2005efficient] present an instance-optimal reconstruction strategy for region families $F$ consisting of points and closed piecewise-$C^{1}$ regions. They do not present a corresponding reconstruction program. When $F$ consists of disks or polygons, a reconstruction program may be derived, albeit with unspecified polynomial-time costs per retrieval. We further discuss their algorithm in the full version. We rely on a key concept of their paper, i.e. a witness.